the perimeter of a rectangle is 76 inches. if the length is two more than five times the width,what is the measurement of the length?
Answer by fractalier(6550)

Perimeter P = 2L + 2W = 76
We also have
L = 5W + 2
Now substitute and get
2(5W + 2) + 2W = 76
10W + 4 + 2W = 76
12W + 4 = 76
12W = 72
W = 6
L = 5(6) + 2 = 32 inches